OpenStudy (anonymous):

Use a graphing calculator to solve the equation -3 cos t = 1 in the interval from 0 to 2p . Round to the nearest hundredth ,please I really need help with this one !!

OpenStudy (anonymous):

I thank you for any help.. I really don't understand this and I have to hand in this worksheet tomorrow am

OpenStudy (anonymous):

If: -3cos(t)=3 then: cos(t)= (-1/3) If you have the cos(angle)=a number , then in order to get the angle, you have to do the inverse cosine function \[\cos^{-1}(the number)=\theta \] here, you'll have \[\cos^{-1}(-1/3)=\theta\]

OpenStudy (anonymous):

the inverse cosine function on your graphic calculator is the same button as cos, but you press 2nd first. Make sure you're in radian mode by pressing the mode button (right next to 2nd), and radian is highlighted in the second row.

OpenStudy (anonymous):

See. they put me in the wrong class and I did not realize I was not this far along..So I don't have a graphing calculator. I tried to find one on-line.. I did ok in pre-algebra but skipped algebra 1 and now I am totally lost but I don't have enough money to transfer to another course

OpenStudy (anonymous):

So I'm super thankful for your help !! is there a way to do this without a graphing calculator?

OpenStudy (anonymous):

yes.... most of the time, the exact answer is left as \(\large \theta = cos^{-1}(-\frac{1}{3}) \). but if they question asked for a decimal approximation, you'd need one...:(

OpenStudy (anonymous):

*the (not they)

OpenStudy (anonymous):

Go here: http://www.wolframalpha.com/ then just type inverse cosine of -1/3. or click this one: http://www.wolframalpha.com/input/?i=inverse+cos+of+-1%2F3&t=sfmtb01

OpenStudy (anonymous):

if you're doing your work at home and are in need of a graphing calculator, this is a very good one... https://www.desmos.com/calculator

OpenStudy (anonymous):

Yes I'm in on-line classes and I have learning issues with math... but because I'm a mom and I can't afford childcare, I have to stick with on-line classses for now.. I just want to set a good example for my kids.. but math is not easy for me

OpenStudy (anonymous):

yes you are setting a good example... good for you..... the people here will gladly help you with any problems you may have... Welcome to OpenStudy...:)

OpenStudy (anonymous):

Thank you so much... So another silly question.. How do I get the alternate signs to come up on the screen/keyboard? like the first part of the answer looks like a funny 0 but with a line going through it.. but I can't find it to type it in my notes..... Is there an alternate way to type it?